OpenStudy (anonymous):

(16x2/3)/(4x1/4) simplify using property of exponents.

OpenStudy (anonymous):

Can you write it in the equation thing...I cant really tell whats going on.

OpenStudy (anonymous):

BTW this website gives a good description of the properties of exponents. http://hotmath.com/hotmath_help/topics/properties-of-exponents.html

OpenStudy (unklerhaukus):

is theis your question \(\dfrac{16x^{2/3}}{4x^{1/4}}\)?

OpenStudy (anonymous):

yes @unckelrhaukus

OpenStudy (unklerhaukus):

use this rule \( \dfrac{x^n}{x^m}=x^{n-m}\)

OpenStudy (anonymous):

thx